# How Reliable Are Automatic Evaluation Methods for Instruction-Tuned LLMs?

### v2

@@ 4 Results and Analysis
In this section, we present a comparative analysis of the evaluation methods in terms of their alignment with human assessments. Our focus lies not on model performance as such, but rather on discerning the reliability and effectiveness of different evaluation methods.
@@ 4.1 Human Evaluation
The ratings of GPT-4 closely aligns with human ratings for correctness, showing slightly more distance in relatedness, and even more in naturalness. (The average differences between human and GPT-4 ratings are summarized in the final row of Table 2.) Based on these results, in the rest of the paper, we focus solely on correctness. We prioritize correctness since it is the most important criterion for determining the usefulness of LLMs.
@@ 4.2 Meta-Evaluation of Metrics
The metric with the highest alignment with human ratings is GPT-4-gold which achieves an average $\mathrm{PA}$ of 0.81 for English short- and long-answer tasks, followed by Rouge-L with 0.75, BertScore with 0.66, and GPT-4-no-gold with 0.62. For all the results see Table 5. Nonetheless, a more fine-grained analysis shows a different pattern, which we will discuss in this section.
Figure 1: Human ratings and Rouge-L scores per task and test language on the bottom and their $\mathrm{PA}$ on top. The human scores are normalized to a range of 0 to 1. Short-answer tasks are on left and long-answer ones on right.
@@
For Tasks 613, 677 and 1659 where the models must generate free-form text, Rouge-L scores are lower than human ratings accompanied with low to medium $\mathrm{PA}$s. This is due to the large set of possible solutions and the small set of gold label answers, often consisting of a single sentence, that the output should match to receive a high Rouge-L score. For some of these tasks, e.g., title generation, it is impossible to cover the set of conceivable solutions which makes Rouge-L unreliable for these types of tasks. The same trend is present for GPT-4, which could be explained by this model’s relying on the gold labels during evaluation, which may make it overly critical of responses that show less conformity with the provided reference answers.
@@
When we remove the gold answers from the prompt, GPT-4’s alignment with human ratings decreases significantly, with $\mathrm{PA}$ dropping from the previous 0.81 to 0.62 for English. Full results are presented in Table 5. For short-answer tasks in English, the reduction is 0.25, and for long-answer tasks, it is 0.09, which underscores GPT-4’s struggle when it lacks a reference for exact matching. We attribute this reduction to the increased complexity of both solving and rating the task, which is more pronounced in Swedish, where there is a higher incidence of models over-generating. While such over-generation may lead to lower human ratings, it could be favored by GPT-4 due to its bias towards longer and more verbose outputs Zheng et al. (2023).
@@
Figure 2: Human and GPT-4-gold’s ratings per task and test language on the bottom and their $\mathrm{PA}$ on top. Ratings are normalized to a range of 0 to 1. Short-answer tasks are on left and long-answer ones on right.
@@
With a $\mathrm{PA}$ at 0.90 for English short-answer tasks, Rouge-L is nearly as well-aligned with human ratings as GPT-4-gold, which has a $\mathrm{PA}$ at 0.93. Our strong results for Rouge-L are in line with findings by Wang et al. (2022), who report a high correlation with humans for classification tasks. With that said, there are instances where Rouge-L has low alignment with human majority rating for short-answer tasks, for example when there is high word overlap between possible answer choices. For example, in Task 1615 the labels are “B entails A”, “B contradicts A”, or “B neutral A”. A wrong answer for this task yields a Rouge-L score of 0.66, as long as the answer is in the possible answers space. The same issue is observed for Task 392 where the label space is “Plausible” and “Not Plausible”.
@@
With a $\mathrm{PA}$ of 0.54 for long-answer tasks, BertScore shows a alignment with humans comparable with GPT-4-gold, which scores 0.58. A comparison of BertScore and GPT-4-gold’s $\mathrm{PA}$s are shown in Figure 3. For complete results of BertScore see Appendix C.3. BertScore achieves comparable results to GPT-4-gold on all long-answer tasks, only underperforming on some of them, particularly Task 1622, where it captures the first criterion which requires a high semantic similarity between the two, but fails to take into account whether enough words have been changed, such as in cases involving synonyms, which is explicitly mentioned in the instructions.
Figure 4: Human and GPT-4-no-gold’s ratings per task and test language on the bottom and their $\mathrm{PA}$ on top. Ratings are normalized to a range of 0 to 1. Short-answer tasks are on left and long-answer ones on right.
@@
For Rouge-L, a reduction of 0.074 in $\mathrm{PA}$ is observed for Swedish compared to English. In contrast, GPT-4-gold experiences no significant reduction when switching from English to Swedish. However, GPT-4-no-gold is less consistent, showing a reduction of 0.044. GPT-4’s decrease in performance when it does not have access to gold references for Swedish outputs suggests that the model has more difficulties solving the task when it is in another language. We believe the effect could be even more pronounced for languages with less training data than Swedish or less typologically similar to English. Further studies would be necessary to investigate this hypothesis.
@@ 5 Conclusions and Future Work
We choose to report our results using pairwise accuracy which we believe provides more robust and reliable alignment statistics compared to common correlation metrics. With that said, $\mathrm{PA}$ has its shortcomings, such as when it faces constant or close-to-constant scores. For example, when the reference vector is $\vec{1}$, a metric vector of $\vec{1}$ or $\vec{3}$ both result in very high $\mathrm{PA}$s, due to the lack of prior knowledge about the metric range. However, this issue is not unique to $\mathrm{PA}$; common correlation metrics also face the same challenge in the case of close-to-constant vectors.
@@ B.1 Translation
Figure 5: Human and BertScore ratings per task and test language on the bottom and their $\mathrm{PA}$ on top. The human ratings are normalized to a range of 0 to 1. Short-answer tasks are on left and long-answer ones on right.
@@ B.2 LLM-as-a-judge
The prompt used for GPT-4 could be found in Table 6. The prompt for GPT-4-no-gold is the same, but without the following part:
@@
Figure 6: The prompt for GPT-4 as evaluator. For GPT-4-no-gold the gold answer is removed.
Appendix C Supplementary Results
C.1 How Good Are Our Language Models at Swedish?
To assess the effectiveness of the models described in Section 3.2, we measure their perplexity. To ensure the generated texts meet high standards and to avoid assessing the models on data used during their pretraining, we use a custom dataset consisting of current news articles from SVT444svt.se, the Swedish national public television broadcaster. The dataset comprises 268 articles spanning various topics, published between June 1st, 2023, and October 16th, 2023. To address the variability caused by different tokenizers across various models, we use character length normalization when calculating perplexity (Liang et al., 2022; Yong et al., 2023).
Model Perplexity
LLaMA2_13b 1.96
LLaMA2_13b_EN 2.03
LLaMA2_13b_SV 2.27
LLaMA2_13b_ENSV 2.24
LLaMA2_7b 2.09
LLaMA2_7b_EN 2.22
LLaMA2_7b_SV 2.51
LLaMA2_7b_ENSV 2.49
SW3_6.7b 1.62
SW3_6.7b_EN 1.65
SW3_6.7b_SV 1.72
SW3_6.7b_ENSV 1.69
Table 3: The perplexity of our models on the SVT dataset. The abbreviations are the training language(s).
C.2 Pairwise Accuracy per Model
Pairwise accuracy per model is shown in Table 4.
Model GPT-4-gold GPT-4-no-gold Rouge-L BertScore
LLAMA2_13b_EN_EN 0.779 0.590 0.601 0.512
LLAMA2_13b_SV_SV 0.816 0.559 0.547 -
LLAMA2_13b_ENSV_EN 0.782 0.594 0.639 0.545
LLAMA2_13b_ENSV_SV 0.828 0.565 0.634 -
LLAMA2_7b_EN_EN 0.766 0.581 0.578 0.454
LLAMA2_7b_SV_SV 0.794 0.578 0.526 -
LLAMA2_7b_ENSV_EN 0.794 0.609 0.607 0.447
LLAMA2_7b_ENSV_SV 0.804 0.577 0.550 -
SW3_6.7b_EN_EN 0.789 0.612 0.604 0.482
SW3_6.7b_SV_SV 0.843 0.645 0.505 -
SW3_6.7b_ENSV_EN 0.830 0.606 0.622 0.491
SW3_6.7b_ENSV_SV 0.828 0.576 0.600 -
Table 4: Pairwise accuracy per model for all metrics.
C.3 BertScore Results
Detailed results of BertScore are shown in Figure 5.
Model Name Task Type Language $\tau$ $\rho$ $\mathrm{PA}$ $\epsilon$
Rouge-L all EN 0.667 0.712 0.752 0.061
Rouge-L long EN 0.308 0.368 0.484 0.061
Rouge-L short EN 0.833 0.871 0.896 0.061
Rouge-L all SV 0.584 0.631 0.699 0.095
Rouge-L long SV 0.325 0.389 0.477 0.095
Rouge-L short SV 0.704 0.743 0.818 0.095
GPT-4-gold all EN 0.781 0.801 0.811 0.000
GPT-4-gold long EN 0.481 0.521 0.584 0.000
GPT-4-gold short EN 0.920 0.930 0.934 0.000
GPT-4-gold all SV 0.792 0.811 0.817 0.000
GPT-4-gold long SV 0.538 0.575 0.601 0.000
GPT-4-gold short SV 0.909 0.920 0.934 0.000
GPT-4-no-gold all EN 0.517 0.535 0.622 0.000
GPT-4-no-gold long EN 0.309 0.338 0.499 0.000
GPT-4-no-gold short EN 0.613 0.625 0.688 0.000
GPT-4-no-gold all SV 0.496 0.514 0.620 0.000
GPT-4-no-gold long SV 0.393 0.425 0.569 0.000
GPT-4-no-gold short SV 0.544 0.556 0.648 0.000
BertScore all EN 0.419 0.482 0.658 0.133
BertScore long EN 0.367 0.452 0.542 0.133
BertScore short EN 0.443 0.495 0.720 0.133
Table 5: Comparison of metrics across different task types for English and Swedish. $\tau$ stands for Kendall’s and $\rho$ for Spearman’s rank correlation coefficient.
Appendix D Implementation Details
Appendix E Human Evaluations
Our evaluators are not crowd-sourced workers; instead, they are individuals with some experience and expertise in the field. They were carefully selected for their familiarity with the subject matter and were hired specifically for this evaluation task. While they were instructed to use their own judgment in the assessment, they also had access to the gold standard answers to guide their evaluations.

### v4

@@ 4 Results and Analysis
In this section, we present a comparative analysis of the evaluation methods in terms of their alignment with human assessments.
@@ 4.1 Human Evaluation
The ratings of GPT-4 closely align with human ratings for correctness, showing slightly more distance in relatedness, and even more in naturalness. (The average differences between human and GPT-4 ratings are summarized in the final row of Table 2.) Based on these results, in the rest of the paper, we focus solely on correctness. We prioritize correctness since it is the most important criterion for determining the usefulness of LLMs. Moreover, comparing our metrics using the other criteria could be problematic. For instance, while Rouge-L scores serve as a reasonable proxy for correctness, they are less suited for evaluating naturalness or relatedness.
@@ 4.2 Meta-Evaluation of Metrics
The metric with the highest alignment with human ratings is GPT-4-gold which achieves an average PA of 0.81 for English short- and long-answer tasks, followed by Rouge-L with 0.75, BertScore with 0.66, and GPT-4-no-gold with 0.62. For a comparison of all the results across different task types, languages, and metrics, see Table A3. Nonetheless, a more fine-grained analysis shows a different pattern, which we will discuss in this section.
Figure 1: Human ratings and Rouge-L scores per task and test language at the bottom and their PA at the top. Human scores are normalized to a range between 0 and 1. Short-answer tasks are on the left and long-answer ones are on the right.
@@
For Tasks 613, 677 and 1659 where the models must generate free-form text, Rouge-L scores are lower than human ratings accompanied with low to medium PAs. This is due to the large set of possible solutions and the small set of gold label answers, often consisting of a single sentence, that the output should match to receive a high Rouge-L score. For some of these tasks, e.g., title generation, it is impossible to cover the set of conceivable solutions which makes Rouge-L unreliable for these types of tasks. The same trend is present for GPT-4, which could be explained by this model’s relying on the gold labels during evaluation, which may make it overly critical of responses that show less conformity with the provided reference answers.
@@
Figure 2: Human and GPT-4-gold’s ratings per task and test language on the bottom and their PA on top. Ratings are normalized to a range of 0 to 1. Short-answer tasks are on left and long-answer ones on right.
@@
When we remove the gold answers from the prompt, GPT-4’s alignment with human ratings decreases significantly, from 0.81 PA to 0.62 PA on English. Full results are presented in Table A3. For short-answer tasks in English, the reduction is 0.25, and for long-answer tasks, it is 0.09, which underscores GPT-4’s struggle when it lacks a reference for exact matching. We attribute this reduction to the increased complexity of both solving and rating the task, which is more pronounced in Swedish, where there is a higher incidence of models over-generating. While such over-generation may lead to lower human ratings, it could be favored by GPT-4 due to its bias towards longer and more verbose outputs Zheng et al. (2023).
@@
With a PA at 0.90 for English short-answer tasks, Rouge-L is nearly as well-aligned with human ratings as GPT-4-gold, which has a PA at 0.93. Our strong results for Rouge-L are in line with findings by Wang et al. (2022), who report a high correlation with humans for classification tasks. With that said, there are instances where Rouge-L has low alignment with human majority rating for short-answer tasks, for example when there is high word overlap between possible answer choices. For example, in Task 1615 the labels are “B entails A”, “B contradicts A”, or “B neutral A”. A wrong answer for this task yields a Rouge-L score of 0.66, as long as the answer is in the possible answers space. The same issue is observed for Task 392 where the label space is “Plausible” and “Not Plausible”.
@@
With a PA of 0.54 for long-answer tasks, BertScore shows a alignment with humans comparable with GPT-4-gold, which scores 0.58. A comparison of BertScore and GPT-4-gold’s PAs are shown in Figure 3. For complete results of BertScore see Appendix C.3. BertScore achieves comparable results to GPT-4-gold on all long-answer tasks, only underperforming on some of them, particularly Task 1622, where it captures the first criterion which requires a high semantic similarity between the two, but fails to take into account whether enough words have been changed, such as in cases involving synonyms, which is explicitly mentioned in the instructions.
Figure 4: Human and GPT-4-no-gold’s ratings per task and test language on the bottom and their PA on top. Ratings are normalized to a range of 0 to 1. Short-answer tasks are on left and long-answer ones on right.
@@
For Rouge-L, a reduction of 0.074 in PA is observed for Swedish compared to English. In contrast, GPT-4-gold experiences no significant reduction when switching from English to Swedish. However, GPT-4-no-gold is less consistent, showing a reduction of 0.044. GPT-4’s decrease in performance when it does not have access to gold references for Swedish outputs suggests that the model has more difficulties solving the task when it is in another language. We believe the effect could be even more pronounced for languages with less training data than Swedish or less typologically similar to English. Further studies would be necessary to investigate this hypothesis.
@@ 5 Conclusions and Future Work
We choose to report our results using pairwise accuracy which we believe provides more robust and reliable alignment statistics compared to common correlation metrics. With that said, PA has its shortcomings, such as when it faces constant or close-to-constant scores. For example, when the reference vector is $\vec{1}$, a metric vector of $\vec{1}$ or $\vec{3}$ both result in very high PAs, due to the lack of prior knowledge about the metric range. However, this issue is not unique to PA; common correlation metrics also face the same challenge in the case of close-to-constant vectors.
@@ Appendix B Prompt Templates
Figure A1: Human and BertScore ratings per task and test language on the bottom and their PA on top. The human ratings are normalized to a range of 0 to 1. Short-answer tasks are on left and long-answer ones on right.
@@ B.2 LLM-as-a-judge
The prompt used for GPT-4 could be found in Table A2. The prompt for GPT-4-no-gold is the same, but without the following part:
@@
Appendix C Supplementary Results
C.1 How Good Are Our Language Models at Swedish?
To assess the effectiveness of the models described in Section 3.2, we measure their perplexity. To ensure the generated texts meet high standards and to avoid assessing the models on data used during their pretraining, we use a custom dataset consisting of current news articles from SVT333svt.se, the Swedish national public television broadcaster. The dataset comprises 268 articles spanning various topics, published between June 1st, 2023, and October 16th, 2023. To address the variability caused by different tokenizers across various models, we use character length normalization when calculating perplexity (Liang et al., 2022; Yong et al., 2023).
Model Perplexity
LLaMA2_13b 1.96
LLaMA2_13b_EN 2.03
LLaMA2_13b_SV 2.27
LLaMA2_13b_ENSV 2.24
LLaMA2_7b 2.09
LLaMA2_7b_EN 2.22
LLaMA2_7b_SV 2.51
LLaMA2_7b_ENSV 2.49
SW3_6.7b 1.62
SW3_6.7b_EN 1.65
SW3_6.7b_SV 1.72
SW3_6.7b_ENSV 1.69
Table A1: The perplexity of our models on the SVT dataset. The abbreviations are the training language(s).
C.2 Pairwise Accuracy per Model
Pairwise accuracy per model is shown in Table A2.
Model GPT-4-gold GPT-4-no-gold Rouge-L BertScore
LLAMA2_13b_EN_EN 0.779 0.590 0.601 0.512
LLAMA2_13b_SV_SV 0.816 0.559 0.547 -
LLAMA2_13b_ENSV_EN 0.782 0.594 0.639 0.545
LLAMA2_13b_ENSV_SV 0.828 0.565 0.634 -
LLAMA2_7b_EN_EN 0.766 0.581 0.578 0.454
LLAMA2_7b_SV_SV 0.794 0.578 0.526 -
LLAMA2_7b_ENSV_EN 0.794 0.609 0.607 0.447
LLAMA2_7b_ENSV_SV 0.804 0.577 0.550 -
SW3_6.7b_EN_EN 0.789 0.612 0.604 0.482
SW3_6.7b_SV_SV 0.843 0.645 0.505 -
SW3_6.7b_ENSV_EN 0.830 0.606 0.622 0.491
SW3_6.7b_ENSV_SV 0.828 0.576 0.600 -
Table A2: Pairwise accuracy per model for all metrics.
C.3 BertScore Results
Detailed results of BertScore are shown in Figure A1.
Model Name Task Type Language $\tau$ $\rho$ PA $\epsilon$
Rouge-L all EN 0.667 0.712 0.752 0.061
Rouge-L long EN 0.308 0.368 0.484 0.061
Rouge-L short EN 0.833 0.871 0.896 0.061
Rouge-L all SV 0.584 0.631 0.699 0.095
Rouge-L long SV 0.325 0.389 0.477 0.095
Rouge-L short SV 0.704 0.743 0.818 0.095
GPT-4-gold all EN 0.781 0.801 0.811 0.000
GPT-4-gold long EN 0.481 0.521 0.584 0.000
GPT-4-gold short EN 0.920 0.930 0.934 0.000
GPT-4-gold all SV 0.792 0.811 0.817 0.000
GPT-4-gold long SV 0.538 0.575 0.601 0.000
GPT-4-gold short SV 0.909 0.920 0.934 0.000
GPT-4-no-gold all EN 0.517 0.535 0.622 0.000
GPT-4-no-gold long EN 0.309 0.338 0.499 0.000
GPT-4-no-gold short EN 0.613 0.625 0.688 0.000
GPT-4-no-gold all SV 0.496 0.514 0.620 0.000
GPT-4-no-gold long SV 0.393 0.425 0.569 0.000
GPT-4-no-gold short SV 0.544 0.556 0.648 0.000
BertScore all EN 0.419 0.482 0.658 0.133
BertScore long EN 0.367 0.452 0.542 0.133
BertScore short EN 0.443 0.495 0.720 0.133
Table A3: Comparison of metrics across different task types for English and Swedish. $\tau$ stands for Kendall’s and $\rho$ for Spearman’s rank correlation coefficient.
Appendix D Implementation Details
Appendix E Human Evaluations
Our evaluators are not crowd-sourced workers; instead, they are individuals with some experience and expertise in the field. They were carefully selected for their familiarity with the subject matter and were hired specifically for this evaluation task. While they were instructed to use their own judgment in the assessment, they also had access to the gold standard answers to guide their evaluations. The instructions given to human evaluators were similar to those given to GPT-4, as presented in Figure 3.
@@ Appendix E Human Evaluations
Figure A2: The prompt for GPT-4 as evaluator. For GPT-4-no-gold the gold answer is removed.